# Dynamic tariffs-based demand response in retail electricity market under uncertainty

### v5

@@ Abstract
Demand response (DR) programs play a crucial role in improving system reliability and mitigating price volatility by altering the core profile of electricity consumption. This paper proposes a game-theoretical model that captures the dynamic interplay between retailers (leaders) and consumers (followers) in a tariffs-based electricity market under uncertainty. The proposed procedure offers theoretical and economic insights by analyzing demand flexibility within a hierarchical decision-making framework. In particular, two main market configurations are examined under uncertainty: i) there exists a retailer that exercises market power over consumers, and ii) the retailer and the consumers participate in a perfect competitive game. The former case is formulated as a mathematical program with equilibrium constraints (MPEC), whereas the latter case is recast as a mixed-integer linear program (MILP). These problems are solved by deriving equivalent tractable reformulations based on the Karush-Kuhn-Tucker (KKT) optimality conditions of each agent’s problem.
Numerical simulations based on real data from the European Energy Exchange platform are used to illustrate the performance of the proposed methodology. The results indicate that the proposed model effectively characterizes the interactions between retailers and flexible consumers in both perfect and imperfect market structures. Under perfect competition, the economic benefits extend not only to consumers but also to overall social welfare. Conversely, in an imperfect market, retailers leverage consumer flexibility to enhance their expected profits, transferring the risk of uncertainty to end-users. Additionally, the degree of consumer flexibility and their valuation of electricity consumption play significant roles in shaping market outcomes. These findings highlight the crucial impact of market structure and consumer behavior on the dynamics of electricity market pricing under demand response programs.
@@
Bi-level programming, Tariff-based demand response, Market power, MILP, MPEC, Retail-consumer equilibrium.
@@ 1 Introduction
[…]
@@
One potential solution is demand response (DR), which is typically achieved through incentive-based or price-based consumer flexibility. These methods offer consumers financial incentives to reduce or shift their peak loads to off-peak periods. To that end, consumers facing a high price during a given hour can adopt one of the following strategies (or a combination of them): i) reducing consumption, ii) shifting consumption to low-price periods, and/or iii) covering part of the consumption by on-site distributive generation or storage sources. This behavior can be used to soften peak loads at given hours and reduce electricity market prices caused by marginal carbon-based technologies (Nolan and O’Malley, 2015; Guo and Weeks, 2022). Moreover, since DR is a distributed resource located at the end of the distribution system, additional environmental benefits are obtained from a reduction of electricity losses in the transmission and distribution lines (Shen et al., 2014). Thus, the main advantages of DR programs are: i) they can contribute to reduce system costs and manage price volatility in real-time (Zugno et al., 2013; Aussel et al., 2020; Alipour et al., 2019), ii) they allow higher penetration of intermittent RES in the electric power system (Morales et al., 2013; Hakimi et al., 2019; Antunes et al., 2020), iii) they help improve system security by contributing to the real-time balance between generation and the demand (Zugno et al., 2013; Wang et al., 2020), and iv) they can mitigate electricity retailers’ market power by increasing consumers flexibility (Devine and Bertsch, 2023).
@@
DR programs employ bidirectional communication to send price signals and financial incentives to responsive consumers, as well as demand signals to retailers or demand aggregators. Upon receiving updated information regarding price tariffs, consumers can calculate their benefits and the trade-off of shifting their consumption at critical times for system operators. The literature on DR is extensive and covers different incentive-based strategies, such as offering consumers dynamic electricity tariffs like time-of-use (TOU), critical peak pricing (CPP), and real-time pricing (RTP). Dynamic tariffs can be attractive to retailers as they allow them to align retail price tariffs with spot market prices, thus transferring some of the retailer-centered risks to customers (Guo and Weeks, 2022; Askeland et al., 2020; Clastres and Khalfallah, 2021). Using tariffs-based DR, Askeland et al. (2020) formulate the electricity grid tariff design problem with a bilevel structure in the context of prosumers at the end-user level. They address the tariff optimization problem as a tool for indirect load control and analyze how a simple tariff scheme can be used to activate end-user flexibility and efficiently reduce grid load by developing Mathematical Problems with Equilibrium Constraints (MPECs). Applying a bi-level programming problem, Clastres and Khalfallah (2021) study the impact of dynamic pricing and load-shifting on a retailer’s electricity supply to consumer markets. Considering the consumer benefit function, Guo and Weeks (2022) model a dynamic price-based DR and find that implementing demand-side management stimuli can reduce consumer electricity demand and further increase retail profit.
@@
However, significant gaps persist in the literature as the integration of RES increases and the demand for consumer flexibility: i) incorporating the characterization of uncertainties affecting electricity prices, production costs, and demands in the day-ahead (DA) market. ii) explicitly studying consumer behavior and leveraging their flexibility for operational and planning decisions. iii) comparing both perfect and imperfect electricity markets, involving multiple participants with varying market powers, to analyze both theoretical and practical economic implications.
@@
Overall, the increase in energy costs, the soaring global competition for fuel, and the tightening of emission requirements place huge pressure on power systems. This motivates us to extend the potential of DR programs to enhance system reliability, ensure supply security, mitigate market power, and manage price volatility. These benefits are achieved while accounting for uncertainty across various market configurations.
@@
The objective of this work is to define meaningful decision-making tools and to analyze potential market interactions between electricity retailers and responsive consumers participating in the day-ahead electricity market. Specifically, we seek to mimic the hierarchy of the decision-making process, where a retailer needs to set first the electricity price tariffs for the consumers, and then they would react accordingly. Moreover, motivated by the concept of smart grids, which promotes different forms of horizontal consumer aggregations (cooperatives, energy communities, virtual power plants, etc.), we assume two market configurations. In the first one, the retailer has market power by anticipating the reaction of consumers, which is modeled by a bilevel (Stackelberg) problem. In the second one, we assume that the retailer is myopic to the consumers’ response, which is modeled with a competitive equilibrium. Furthermore, for both cases and extending previous works, we improve the characterization of the consumers’ demand response by using a (non-linear) quadratic utility function.
@@
These facts, together with accounting for different levels of market competition and sources of uncertainty, have a great impact on the retailers’ and consumption decision-making processes, and ultimately on the resulting market outcomes. Therefore, this work models the profit maximization problems of retailers and the utility maximization of consumers in electricity consumption, taking into account the inherent uncertainties present in both perfect and imperfect electricity markets. Specifically, the retailer’s problem with market power is formulated as a bi-level program and solved as an MPEC (the interested reader is referred to Luo et al. (1996) for a complete treatment of bilevel programs and their solution methods), and the equilibrium model with perfect competition is solved as a mixed-integer linear programming (MILP) (the interested reader is referred to Leyffer and Munson (2010) for a complete treatment of how equilibrium problems can be solved as mixed complementarity problems (MCP)). Uncertainty is dealt with stochastic programming techniques, which extends the majority of deterministic DR models available in the literature.
@@
The paper presents several key findings: i) it demonstrates that retailers can maximize their expected profits by exercising market power and utilizing consumer flexibility to mitigate risk. ii) it shows that consumers can minimize their procurement costs, and increase overall social welfare under a perfect equilibrium model. iii) the proposed models are versatile and capable of adapting to various groups of consumers with flexible demand across different market configurations. iv) explicitly modeling consumer behavior through a utility function in a dynamic pricing environment provides deeper insights into characterizing end-users. This approach reveals how their flexibility can aid the integration of renewable energy sources into the power system under DR programs. A key takeaway is the increased vulnerability of the electricity market to the market power of oligopolistic entities when demand-side flexibility is neglected, leading to market inefficiencies. However, harnessing consumer flexibility in operational and strategic planning decisions can counterbalance retailers’ market power and mitigate grid instability.
@@
The remainder of this paper is organized as follows. Section 2 reviews the current literature on the topic. Section 3 demonstrates the model formulation which entails the retailer’s and consumers’ problems. The formulation of the problems as an MPEC, and linear reformulation of the problem as MILP in the equilibrium are presented. Computational and simulation results are reported and discussed in Section 4. Finally, Section 5 closes by drawing some concluding remarks from the provided discussions and results.
@@ 2 Literature review
From the supply side, power producers, encompassing both conventional and RES, sell their generated power to the grid through organized energy markets. These markets are responsible for handling transactions related to energy and ancillary services. On the demand side, DR programs play an integral role. These programs, managed by DR buyers such as load-serving entities, distribution system operators (DSOs), large consumers, and DR aggregators, are pivotal in implementing strategies to collect and manage flexibility options from end-users. This management effectively modifies energy consumption patterns in response to supply conditions or market signals.
@@
DR buyers, while engaging in various market activities, primarily focus on interacting with consumers through contracts or DR programs to create system flexibility. Such flexibility is particularly crucial in day-ahead and intraday markets, where supply and demand adjustments occur closer to real-time operations. The specific nature and extent of these interactions and transactions can vary, depending on the particular market design and regulatory environment.
@@
In the literature, DR programs have been considered pivotal in addressing a range of economic and operational challenges. These challenges include achieving economic efficiency, reducing market volatility, optimizing utility in electricity consumption, minimizing discomfort in existing consumption patterns, enhancing grid stability, optimizing system performance, and reducing the carbon footprint. Studies such as those by Devine and Bertsch (2023); Guo and Weeks (2022); Askeland et al. (2020); Clastres and Khalfallah (2021); Nilsson et al. (2018), among others have extensively explored these facets. From a retail perspective, the importance of dynamic tariffs-based DR is particularly noteworthy. These tariffs align retail prices with spot market prices, thereby transferring some of the risks centered around retailers to the customers. As highlighted by Nilsson et al. (2018), dynamic tariffs are recognized as one of the most effective strategies to enhance demand flexibility. This strategy not only complements the DR programs but also serves as a critical link between fluctuating market prices and consumer behavior.
Furthermore, the role of DR in general, and specifically in mitigating market power, is examined by Devine and Bertsch (2023) using a market equilibrium model. This study underscores the importance of distinguishing between regular consumers and prosumers with distributed energy resources (DERs). Such a distinction is crucial for a quantitative analysis of the effects and benefits of DR programs, illustrating the nuanced ways in which DR programs can influence market dynamics and consumer participation.
@@
The demand response literature, from a theoretical perspective, has made significant strides in addressing market complexities. It encompasses studies considering imperfect markets using bilevel models, such as Afşar et al. (2016); Soares et al. (2019); Aussel et al. (2020); Soares et al. (2020) and perfect competition with equilibrium models such as Devine and Bertsch (2023, 2018); Hu et al. (2016); Nguyen et al. (2016).
@@
Afşar et al. (2016) present a bi-level optimization model where the supplier establishes time-differentiated electricity prices, and the consumers minimize their electricity consumption costs to enhance grid efficiency. Soares et al. (2019) present a model considering a single leader (retailer) whose aim is to maximize profit and multi-followers (consumers) whose objective is to minimize their electricity cost. Aussel et al. (2020) propose the interactions among electricity suppliers, local agents, aggregators, and consumers with a trilevel multi-leader-multi-follower game for load shifting induced by the time of use pricing. Soares et al. (2020) develop a comprehensive model in which a retailer maximizes its profit and a cluster of consumers react to the decision of the retailer based on electricity prices by determining the operation of the controllable loads to minimize the electricity bill and a monetized discomfort factor associated with the indoor temperature deviations.
@@
Moreover, DR programs have been extensively analyzed using equilibrium models in various electricity market applications. The widespread application of these models in electricity market research stems from their robust ability to represent diverse market structures and the interactions between market participants. For instance, Devine and Bertsch (2023) investigate the role of DR in mitigating market power, particularly through load shifting and self-generation. Devine and Bertsch (2018) delves into the costs associated with providing flexibility through load shedding and self-generation using DERs. Nguyen et al. (2016) analyzes the energy scheduling challenges faced by load-serving entities, particularly in managing flexible and inflexible loads. Additionally, Hu et al. (2016) present a stochastic-multi-objective Nash-Cournot model, which aims to reduce peak demand, energy prices, and emissions.
@@
Furthermore, the application of tariff-based DR strategies has been explored in various contexts. For distribution network (Lu et al., 2018), optimal bidding strategy in day-ahead electricity markets (Vayá and Andersson, 2014), optimal pricing strategy in pool-based electricity markets (Ruiz and Conejo, 2009), modeling coordinated cyber-physical attacks (Li et al., 2015), and optimal charging schedules of plug-in electric vehicles (Momber et al., 2015), among others. Summarizing the benefits of DR programs from the perspective of various electricity market participants is a complex task, given the extensive range of literature on the subject. However, these benefits have been summarized in Figure 1.
@@
Figure 1: Enhanced summary of DR programs benefits.
The existing literature on dynamic electricity price tariffs extensively examines their impact on individual players in the energy market. This includes areas such as cost minimization for demand aggregators, optimal reconfiguration of microgrids by DSOs, optimal allocation of distributed generations by TSOs, retailer profit maximization, and consumer benefit enhancement. Notable studies in this domain include those by Ajoulabadi et al. (2020); Wang et al. (2020); Nejad et al. (2019); Dagoumas and Polemis (2017); Nilsson et al. (2018), among others. The task of explicitly modeling consumer behavior in response to dynamic pricing is particularly challenging in this field. This entails understanding their price elasticities or utility functions, particularly during uncertain times. Such modeling remains a demanding and highly complex research topic in DR programs. Addressing different market configurations, including both strategic and myopic retailers, and explicitly modeling consumer behavior through utility functions further complicates this task. The integration of stochastic and non-dispatchable RES makes it difficult to manage these systems efficiently.
By leveraging the advantages of game-theoretical models to address market power and competitive markets, in this work, we develop a tariffs-based demand response program by simultaneously solving retailers’ and consumers’ problems. Assuming that consumers are rational decision-makers and risk-averse to their electricity consumption, we explicitly model their behavior, considering utility functions to quantify their welfare. While measuring satisfaction is an abstract concept, in classical microeconomics, individuals’ behavior is analyzed by the choices they make when purchasing a product, which is represented by utility functions or indifference curves (Rubinstein et al., 2006).
Thus, the main contributions of this paper are:
@@
    Application of demand response to address retailer and consumer decision-making challenges in electricity markets with dynamic pricing. Imperfect competition is analyzed using a Stackelberg game and perfect competition is through equilibrium models, incorporating uncertainties in RES and demand. The focus is on optimizing social welfare and mitigating market power through the strategic use of consumer flexibility.
@@
    The retailer-consumer market power interaction is transformed into a single-level problem using a tractable MPEC. On the other hand, the equilibrium is addressed as an MILP problem, where consumer demand is treated as a variable for both upper- and lower-level problems, and dynamic pricing is considered an exogenous parameter. In addition, an NLP reformulation is employed, providing solutions that are equivalent to those of the MILP for the equilibrium problem.
@@
    Introducing a utility function to quantify consumer benefits from flexibility, which, together with the modeling of battery storage, extends consumers’ flexibility to two possibilities: i) reacting to the price tariff where they can decide to buy more or less energy at a particular hour, and ii) delaying or anticipating consumption within hours. Modeling these two flexibilities has significant operational and economic implications.
@@
    Utilizing an extensive set of real-world case studies to illustrate the effectiveness of our model in capturing market structures and the implications of DR programs on the operational and economic decisions of market participants. The numerical results focus on analyzing the impact of consumer flexibility on welfare and other market outcomes. The insights derived from this analysis can be useful, not only for market participants and regulators but also to guide future research in this field.

### v6

# Dynamic tariff-based demand response in retail electricity market under uncertainty

@@ Abstract
Demand response (DR) programs play a crucial role in improving system reliability and mitigating price volatility by altering the core profile of electricity consumption. This paper proposes a game-theoretical model that captures the dynamic interplay between retailers (leaders) and consumers (followers) in a tariff-based electricity market under uncertainty. The proposed procedure offers theoretical and economic insights by analyzing consumer flexibility within a hierarchical decision-making framework. In particular, two main market configurations are examined under uncertainty: i) there exists a retailer that exercises market power over consumers, and ii) the retailer and the consumers participate in a perfect competitive game. The former is formulated as a mathematical program with equilibrium constraints (MPEC), whereas the latter is recast as a mixed-integer linear program (MILP). These problems are solved by deriving equivalent tractable reformulations based on the Karush-Kuhn-Tucker (KKT) optimality conditions of each agent’s problem.
Numerical simulations based on real data from the European Energy Exchange platform are used to illustrate the performance of the proposed methodology. The results indicate that the proposed model effectively characterizes the interactions between retailers and flexible consumers in both perfect and imperfect market structures. Under perfect competition, the economic benefits extend not only to consumers but also to overall social welfare. Conversely, in an imperfect market, retailers leverage consumer flexibility to enhance their expected profits, transferring the risk of uncertainty to end-users. Additionally, the degree of consumer flexibility and consumers’ valuation of electricity consumption play significant roles in shaping market outcomes. These findings highlight the crucial impact of market structure and consumer behavior on the dynamics of electricity market pricing under demand response programs.
@@
Bilevel programming, Consumer flexibility, Tariff-based demand response, Market power, Retail-consumer equilibrium, Utility function.
@@ 1 Introduction
The integration of renewable energy sources (RES) is rapidly increasing, driven by their environmentally friendly properties and economic benefits. Key global players such as the European Union, the United States, and China have set ambitious RES generation targets, aiming for 100%, 80%, and 60%, respectively, by the year 2050. This green transition aims at reducing the reliance on emitting power sources, thereby decreasing CO2 emissions and advancing towards achieving net-zero carbon goals. Despite these benefits, high RES penetration introduces significant uncertainty and system imbalances. These factors present considerable challenges to maintaining the reliability and security of power systems. Hence, it is of key importance to determine optimal operation strategies in electricity markets that ensure the real-time balance of supply and demand, considering the high penetration of RES and demand uncertainties (He et al., 2013; Zugno et al., 2013; Niromandfam et al., 2020; Abate et al., 2022). For instance, most of the markets where electricity is traded are contracted and cleared weeks, days, or hours before the energy delivery takes place and uncertainty is resolved. In such cases, the market may impose penalties or incur high curtailment costs if the dispatched power is not consumed as planned in real-time. Moreover, this imbalance can compromise system stability and requires the use of fast-response conventional units. Therefore, to enable the large-scale integration of RES and to enhance the decarbonization of electricity systems without endangering the security of supply, additional flexibility must be provided in the form of demand-side flexibility (Wang et al., 2020; Morales-España et al., 2021).
@@
One potential solution is demand response (DR), which is typically achieved through incentive-based or price-based consumer flexibility. It works by offering consumers financial incentives to reduce or shift their peak loads to off-peak periods. To that end, consumers facing a high price during a given hour can adopt one of the following strategies (or a combination of them): i) reducing consumption, ii) shifting consumption to low-price periods, and/or iii) covering part of the consumption by on-site distributive generation or storage sources. This behavior can be used to soften peak loads at given hours and reduce electricity market prices caused by marginal carbon-based technologies (Nolan and O’Malley, 2015; Guo and Weeks, 2022). Moreover, since DR is a distributed resource located at the end of the distribution system, additional environmental benefits are obtained from a reduction in electricity losses in the transmission and distribution lines (Shen et al., 2014). Thus, the main advantages of DR programs are: i) they can contribute to reducing system costs and managing price volatility in real-time (Zugno et al., 2013; Aussel et al., 2020; Alipour et al., 2019); ii) they allow higher penetration of intermittent RES in the electric power system (Morales et al., 2013; Hakimi et al., 2019; Antunes et al., 2020); iii) they help improve system security by contributing to the real-time balance between generation and demand (Zugno et al., 2013; Wang et al., 2020); and iv) they can mitigate electricity retailers’ market power by increasing consumer flexibility (Devine and Bertsch, 2023).
@@
Upon receiving updated information regarding price tariffs, consumers can calculate their benefits and the trade-off of shifting their consumption at critical times for system operators. The literature on DR extensively explores various incentive-based strategies. These include offering consumers dynamic electricity tariffs, such as time-of-use (TOU), critical peak pricing (CPP), and real-time pricing (RTP). Dynamic tariffs can be attractive to retailers because they allow them to align retail price tariffs with spot market prices, thus transferring some of the retailer-centered risks to customers (Guo and Weeks, 2022; Askeland et al., 2020; Clastres and Khalfallah, 2021). Using tariff-based DR, Askeland et al. (2020) formulate an electricity grid tariff design problem with bilevel programming in the context of prosumers at the end-user level. They address the tariff optimization problem as a tool for indirect load control. By developing mathematical problems with equilibrium constraints (MPECs), they analyze how a simple tariff scheme can enhance prosumers’ flexibility and efficiently reduce grid load. Applying a bilevel programming problem, Clastres and Khalfallah (2021) study the impact of dynamic pricing and load-shifting on a retailer’s electricity supply to consumer markets. In a related study, Guo and Weeks (2022) model a dynamic price-based DR program, incorporating the consumer benefit function. They find that implementing demand-side management incentives can reduce consumer electricity demand during peak hours and increase retailer profits.
@@
However, significant gaps persist in the literature as the integration of RES grows and the demand for consumer flexibility increases. Key areas include: i) incorporating the characterization of uncertainties affecting electricity prices, production costs, and demand in the day-ahead market; ii) explicitly studying consumer behavior and leveraging flexibility for operational and planning decisions; and iii) comparing perfect and imperfect electricity markets, involving multiple participants with varying market powers to analyze both theoretical and practical economic implications.
@@
Overall, the increase in energy costs, the global competition for fuel, and the tightening of emission requirements place considerable pressure on power systems. This motivates us to extend the potential of DR programs to enhance system reliability, ensure supply security, mitigate market power, and manage price volatility. These benefits are achieved while accounting for uncertainty across various market configurations.
@@
The objective of this work is to model and analyze the decision-making process and potential market interactions between electricity retailers and flexible consumers participating in the day-ahead electricity market. Specifically, we seek to mimic the hierarchy of the decision-making process, where a retailer needs to first set electricity price tariffs and the consumers react accordingly. Moreover, motivated by the concept of smart grids, smart devices, such as batteries, EVs, or heat pumps, which promote different forms of horizontal consumer aggregations (cooperatives, energy communities, virtual power plants, etc.), we assume two market configurations. In the first one, the retailer has market power by anticipating the reactions of consumers, which is modeled by a bilevel (Stackelberg) problem. In the second one, we assume that the retailer is myopic to the consumers’ response, which is modeled with a competitive equilibrium. Furthermore, for both configurations and extending previous works, we model the characterization of the consumers’ flexibility by using a (nonlinear) quadratic utility function.
@@
These factors, along with varying levels of market competition and sources of uncertainty, influence retailers’ and consumers’ decision-making processes. This, in turn, impacts the resulting market outcomes. Therefore, this work models the profit maximization problems of retailers and the utility maximization problems of consumers in electricity consumption, taking into account the inherent uncertainties present in both perfect and imperfect electricity markets. Specifically, the retailer’s problem with market power is formulated as a bilevel program and solved as an MPEC. For a comprehensive treatment of bilevel programs and their solution methods, the interested reader is referred to Luo et al. (1996). The equilibrium model with perfect competition, on the other hand, is solved via mixed-integer linear programming (MILP). We model uncertainty with stochastic programming, which extends the majority of deterministic DR models available in the literature.
@@
This paper presents several key findings: i) It demonstrates that retailers can maximize their expected profits by exercising market power and utilizing consumer flexibility to mitigate/transfer their risk. ii) The results show that consumers can minimize their electricity consumption costs and increase overall social welfare under a perfectly competitive market. iii) The proposed models can be adapted to various groups of consumers with flexibility. They can also include prosumers who trade electricity, either jointly or in equilibrium, to achieve individual objectives across different market configurations. iv) Explicitly modeling consumer behavior through a utility function in a dynamic pricing environment offers managerial insights. This approach helps to characterize and leverage consumer flexibility to optimize social welfare, especially with high RES penetration. Another takeaway is the increased vulnerability of the electricity market to the market power of oligopolistic entities when consumer flexibility is neglected, leading to market inefficiencies. Harnessing consumer flexibility in operational and strategic planning decisions with a competitive market structure can counterbalance retailers’ market power and mitigate grid instability.
@@
The remainder of this paper is organized as follows. Section 2 reviews the current literature on this topic. Section 3 demonstrates the model formulation, which entails the retailer’s and consumers’ problems. The formulation of the problems as an MPEC and the linear reformulation of the problem as an MILP in equilibrium is presented. The computational and simulation results are reported and discussed in Section 4. Finally, Section 5 closes by drawing some concluding remarks from the provided discussions and results.
@@ 2 Literature review
On the supply side, power producers, both conventional and RES, sell their generated power to the grid through organized energy markets. These markets are responsible for handling transactions related to energy and ancillary services. On the demand side, DR programs play an integral role in balancing and operational security in power systems. These programs, managed by DR buyers such as load-serving entities, distribution system operators (DSOs), large consumers, and DR aggregators, are pivotal in implementing strategies to collect and manage flexibility options from end-users. This management effectively modifies energy consumption patterns in response to supply conditions or market signals.
@@
DR buyers primarily focus on interacting with consumers through contracts to create system flexibility. DR programs are crucial in day-ahead and intra-day markets, where supply and demand adjustments occur closer to real-time operations. The specific nature and extent of these interactions and transactions can vary, depending on the particular market design and regulatory environment.
@@
In the literature, DR programs have been considered to address various economic and operational challenges. These challenges include achieving economic efficiency, reducing market volatility, optimizing utility in electricity consumption, minimizing discomfort in existing consumption patterns, enhancing grid stability, optimizing system performance, and reducing the carbon footprint. Devine and Bertsch (2023), Guo and Weeks (2022), Askeland et al. (2020), Clastres and Khalfallah (2021) andNilsson et al. (2018), among others, have explored and proposed models to address some of these challenges. From retailers’ perspective, dynamic tariff-based DR can be used in spot markets to transfer some of retailers’ risks to consumers. As highlighted by Nilsson et al. (2018), dynamic tariffs are recognized as one of the most effective strategies for enhancing demand flexibility. This strategy not only complements DR programs but also serves as a critical link between fluctuating market prices and consumer behavior.
The role of DR in mitigating market power is examined by Devine and Bertsch (2023) using a market equilibrium model. They underscore the importance of distinguishing between regular consumers and prosumers with distributed energy resources (DERs). Such a distinction is crucial for a quantitative analysis of the effects and benefits of DR programs, illustrating the nuanced ways in which DR programs can influence market dynamics and consumer participation. For instance, Jens et al. (2022) discuss approaches for leveraging prosumers’ flexibility as balancing service providers in power systems, either through joint clustering or equilibrium settings. Prosumers can trade with other prosumers (peer-to-peer trading (P2P)) or act as a larger unit within traditional electricity markets. One approach involves jointly clustering prosumers for a common goal within the microgrid, leveraging batteries, EVs, and heat pumps without explicitly modeling these devices. Another approach allows prosumers to act selfishly and exploit their flexibility in equilibrium, from which no participant can improve by unilaterally deviating from the equilibrium.
@@
From a theoretical perspective, the demand response literature has made significant strides in addressing market complexities. It encompasses studies considering imperfect markets using bilevel models, such as Afşar et al. (2016), Soares et al. (2019), Aussel et al. (2020), and Soares et al. (2020) and perfect competition with equilibrium models, such as Devine and Bertsch (2023, 2018); Hu et al. (2016); Nguyen et al. (2016).
@@
Afşar et al. (2016) present a bilevel optimization model in which the supplier establishes time-differentiated electricity prices and consumers minimize their electricity consumption costs to enhance grid efficiency. Soares et al. (2019) present a model considering a single leader (retailer) whose aim is to maximize profit and multiple followers (consumers) whose objective is to minimize electricity costs. Aussel et al. (2020) propose interactions among electricity suppliers, local agents, demand aggregators, and consumers with a tri-level multileader-multi follower game for load shifting induced by the time of use pricing. Soares et al. (2020) develop a comprehensive model in which a retailer maximizes its profit. A cluster of consumers responds to the retailer’s decisions on electricity prices by adjusting the operation of controllable loads to minimize their electricity bills and monetize a discomfort factor based on indoor temperature deviations.
@@
Moreover, DR programs have been extensively analyzed using equilibrium models in various electricity market applications. The widespread application of these models in electricity market research stems from their robust ability to represent diverse market structures and interactions between market participants. For instance, Devine and Bertsch (2023) investigate the role of DR in mitigating market power, particularly through load shifting and self-generation. Devine and Bertsch (2018) delve into the costs associated with providing flexibility through load shedding and self-generation using DERs. Nguyen et al. (2016) analyze the energy scheduling challenges faced by load-serving entities in managing flexible and inflexible loads. Hu et al. (2016) present a stochastic-multi-objective Nash-Cournot model to reduce peak demand, energy prices, and emissions.
@@
Furthermore, the application of tariff-based DR strategies has been explored to address other various challenges. These include distribution networks (Lu et al., 2018), optimal bidding strategies in day-ahead electricity markets (Vayá and Andersson, 2014), optimal pricing strategies in pool-based electricity markets (Ruiz and Conejo, 2009), modeling coordinated cyber-physical attacks (Li et al., 2015), and optimal charging schedules for plug-in electric vehicles (Momber et al., 2015), among others. Summarizing the benefits of DR programs from the perspective of various electricity market participants is a complex task, given the extensive range of literature on the subject. Some of the benefits are summarized in Figure 1.
@@
Figure 1: Enhanced summary of DR program benefits.
The literature on dynamic electricity price tariffs extensively examines their impact on individual players in the energy market. This includes cost minimization for demand aggregators, optimal reconfiguration of microgrids by DSOs, optimal allocation of distributed generations by TSOs, retailer profit maximization, and consumer benefit enhancement (Ajoulabadi et al., 2020; Wang et al., 2020; Nejad et al., 2019; Dagoumas and Polemis, 2017; Nilsson et al., 2018).
By leveraging the advantages of game-theoretical models to address market power and competitive markets, in this work, we develop a tariff-based demand response program by simultaneously solving retailers’ and consumers’ problems. Assuming that consumers are rational decision makers and risk averse to their electricity consumption, we explicitly model their behavior by considering utility functions to quantify their welfare. Explicitly modeling human behavior is challenging, particularly in practical implementations, let alone determining its functional form. However, using utility functions to measure satisfaction is common in the classical literature, such as in performance evaluations of scheduling algorithms for telecommunication systems, where utility functions are widely used (Palomar and Chiang, 2007). The specific choice of a utility function to measure such behavior usually depends on the nature of the problem, the pattern of benefits under consideration, and the required mathematical properties, among other factors. In demand response programs, utility functions can i) be used as the statistical expectation of customers’ evaluation of electricity consumption (Samadi et al., 2012); ii) illustrate how a rational consumer would make consumption decisions considering incentives from reducing/shifting electricity; and iii) in practical settings, be inferred/imposed by tariffs under different market structures, such as those involving myopic and strategic retailers (Jens et al., 2022; Niromandfam et al., 2020). Specifically, in settings where we can assume that consumers do not frequently switch to a different retailer (short-term model), the utility function should capture a decrease in its marginal value with the consumption level (diminishing returns). This could translate into demand curves where consumption levels decrease monotonically with rising prices, as expected for rational consumers. Such demand behavior would also prevent retailers from excessively increasing price tariffs to maximize profit (Zugno et al., 2013). Thus, the main contributions of this paper are:
@@
    Applying demand response to address retailers’ and consumers’ decision-making challenges in electricity markets with dynamic pricing. Imperfect competition is analyzed using a Stackelberg game, and perfect competition is analyzed through equilibrium models, accounting for uncertainties by employing stochastic programming.
@@
    Tackling the bilevel problem between the retailer and consumers by directly solving the MPEC formulation without linearization, leveraging NLP solvers to handle the complexity of complementarity constraints. In our equilibrium model, consumer demand is treated as a variable in both upper- and lower-level problems, and dynamic prices are considered as exogenous parameters. The equilibrium value will naturally come out as the market clearing price that satisfies all the players’ (retailers’ and consumers’) optimality conditions. This approach contrasts with typical bilevel models, where price tariffs are treated as endogenous when transforming to MILP.
@@
    Introducing a utility function to quantify consumer flexibility, which, together with the availability of battery storage and other smart devices, extends consumer flexibility to two possibilities: i) reacting to the price tariff where they can decide to buy more or less energy at a particular hour, and ii) delaying or anticipating consumption within hours.
@@
    Utilizing an extensive set of real-world case studies to illustrate the effectiveness of our model in capturing market structures and the implications of DR programs for the operational and economic decisions of market participants. The numerical results focus on analyzing the impact of consumer flexibility on social welfare and other market outcomes.